OpenStudy (anonymous):

Solve: 2x + 3y + z = -3 x + 4y - 3z = -23 3x - y +2z = 14 I don't even know how to start this problem:(

OpenStudy (anonymous):

have u learnt Cramer's rule?

OpenStudy (anonymous):

or the matrix method?

OpenStudy (anonymous):

No I have not. At least I don't believe so. I know I need to reduce the problem down to like 2 equations

OpenStudy (anonymous):

then which method would u adopt for the solution?

OpenStudy (anonymous):

we can reduce the equations into two variables n two equtations

OpenStudy (anonymous):

would I need to set the equations equal to zero and factor?

OpenStudy (anonymous):

no u have to eliminate one variable, then solving them simultaneously

OpenStudy (anonymous):

like, we can choose eq 1 and 2 n eliminate x,say

OpenStudy (anonymous):

likewise we can select eq 2 n three, again eliminate x

OpenStudy (anonymous):

we get a et of two eqs in 2 variables

OpenStudy (anonymous):

got it?or u need the solution?

OpenStudy (anonymous):

Can you show me the steps? That would help me a lot.

OpenStudy (anonymous):

I am studying for my final and this is a practice problem I am doing, I just can't figure it out.

OpenStudy (anonymous):

multiply eq 2 with 2 and subtract eq 1 and 2 2x+3y+z=-3 2x-2y+4z=28 ---------------------- 5y-3z=-31(on subtraction)

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

similarly multiplying eq 2 by 3 and subtracting 2 and 3

OpenStudy (anonymous):

3x+12y-9z=-69 3x-y+2z=14 --------------- -13y-11z=-83

OpenStudy (anonymous):

is it ok?

OpenStudy (anonymous):

yes, Thank you.

OpenStudy (anonymous):

now there r two equations in two variables u can use either substitution method or elimination method

OpenStudy (anonymous):

2x+3y+z=-3 2x+8y-6z=-46 --------------- -5y+7z=-43

OpenStudy (anonymous):

i missed the multiplication by 2 first :(

OpenStudy (anonymous):

I don't know, its just still very confusing to me. So can you go back to the beginning..... I have 2x + 3y + z = -3 but where did the equation blow it come from?